# MeanSparse: Post-Training Robustness Enhancement Through Mean-Centered Feature Sparsification

### v1

@@ 4.3 Insights from Ablation Study
We conducted several experiments to explore different aspects of MeanSparse integration into trained models. These experiments examine the effects of activation functions A.1, the type of adversarial attack (either $\ell_{\infty}$ or $\ell_{2}$) A.2, the types of adversarial examples used during training with AT A.3, the impact of centering MeanSparse around the mean value A.4, the influence of attack power on MeanSparse integration performance A.5, and the effects of integrating MeanSparse with different patterns, such as isolated integration with a single activation A.6. The main takeaways from our ablation study can be summarized as:
@@
  • Although we have several results showing performance improvements in convolutional neural networks (CNNs), the integration of MeanSparse into attention-based architectures is still under consideration. In our initial experiment, we have not observed results on par with those in CNN-based architectures.
@@ 4.4 Limitations
The MeanSparse technique proposed in this paper is a post-training method that is useful on top of adversarially trained models. Using MeanSparse for models trained in a standard way does not lead to robustness improvement as we expect based on our intuition introduced in this paper. The current version of the MeanSparse technique has been applied to convolutional neural networks, and we still have not achieved comparable results on attention-based architectures. The current work explores the performance of the MeanSparse technique on the CIFAR-10 and ImageNet-1k datasets, which can be extended in future works.
@@ 5 Conclusions
Adversarial training, a widely accepted defense against evasion attacks, attenuates non-robust feature extractors in a model. Although adversarial training (AT) is effective, this paper demonstrates that there is still an easily accessible capacity for attackers to exploit. We introduce the MeanSparse technique, which partially blocks this capacity in the model available to attackers and enhances adversarial robustness. This technique, easily integrable after training, improves robustness without compromising clean accuracy. By integrating this technique into trained models, we achieve a new record in adversarial robustness against $\ell_{\infty}$ untargeted attacks on both CIFAR-10 and ImageNet-1k datasets.
@@ Appendix A Ablation Experiment
We conducted several experiments to evaluate the effectiveness of MeanSparse in various scenarios. In this section, we use APGD accuracy (accuracy after attacking the model with Auto-PGD using both Cross-entropy and Difference of Logits Ratio loss metrics) as the robustness metric [8]. During our simulations, we found APGD accuracy to be almost equal to AA accuracy, with the advantage that APGD can be measured faster.
@@ A.1 Activation Function
Table 2: Clean and APGD accuracy of ResNet18 model with different activation functions over CIFAR-10 test set before (Base) and after integration with MeanSparse technique for different value of $\alpha$
@@ A.2 Threat Model
In this experiment, we focused on the effectiveness of MeanSparse against two well-studied threat models, $\ell_{\infty}$ and $\ell_{2}$. Until now, our focus was on the $\ell_{\infty}$ threat model, which generally results in lower robust accuracy [7]. On the other hand, $\ell_{2}$ threat models can focus on a specific part of the image, and this localization of adversarial perturbation may challenge the effectiveness of the MeanSparse technique. In this experiment, we compare the robustness against both $\ell_{\infty}$ and $\ell_{2}$ threat models. The model is ResNet18 with the GELU activation function as in our Activation Function experiment in 2. Table 3 illustrates the clean and APGD accuracy for $\ell_{\infty}$ and $\ell_{2}$ threats. As the accuracy for the $\ell_{\infty}$ threat model increases with increasing $\alpha$, the accuracy for $\ell_{2}$ also increases, which depicts the effectiveness and generalization of MeanSparse across well-studied threat models. Additionally, the improvement compared to the base model at $\alpha=0.2$ is larger for the $\ell_{2}$ threat than for the $\ell_{\infty}$ threat.
Table 3: Clean and APGD accuracy of ResNet18 model with GELU activation function over CIFAR-10 test set before (Base) and after integration with MeanSparse technique for different value of $\alpha$ and different threat models from RobustBesnch [8]
@@
A.3 Adversarial Training
@@
A.4 Feature Centerization Type
The MeanSparse technique applies sparsification over mean-centered features across each channel in the input tensor to the activation function. In this section, we inspect the importance of the reference for feature centralization before sparsification in our proposed technique. For this purpose, we compare three cases where the reference for centering features is the channel mean ($\boldsymbol{\mu}_{ch}$), zero, and the mean over all features in all channels ($\mu$). For a fair comparison across all cases, we select the sparsification threshold equal to $\alpha\times\sigma_{ch}$, where $\sigma_{ch}$ emphasizes the channel-wise calculation standard deviation. Table 5 represents the result for different options of reference. When using zero as the reference for feature centralization, the improvement in APGD is higher than using the mean as the reference, but on the other hand, the utilization of the model also decreases. One challenging problem is the high rate of clean accuracy reduction when using zero as the reference, which may cause a huge reduction if not selected properly. When using the mean of all channels as the reference, the results in terms of both clean and APGD accuracy metrics underperform compared to the case where the channel mean is used as the reference. The main takeaway from the table is the importance of using channel-wise statistics to maintain utility while improving robustness.
Table 5: Clean and APGD accuracy of ResNet18 model with GELU activation function over CIFAR-10 test set before (Base) and after the application of MeanSparse technique for different references for feature centralization
@@ A.5 Attack Power
Table 6: APGD accuracy of ResNet18 model with GELU activation function over CIFAR-10 test set before (Base denoted by $\alpha=0$ in the table) and after the application of MeanSparse technique for different attack power $\epsilon$ and two different $\alpha$ values ($0.20$ and $0.35$). Clean accuracy is provided next to each $\alpha$ value.
@@
A.6 Applying to Specific Activation Functions
The MeanSparse technique is designed to sparsify mean-centered inputs to all activation functions in a model. In this experiment, we will focus on applying sparsification over a selected subset of activation functions. For this purpose, we index the activation functions in the ResNet18 model. This model has 8 residual blocks, each with 2 activation functions: one in the Main Path and one after the result of the main path is added to the residual connection, totaling 16 activation functions. Additionally, this architecture has one activation function before the residual block. So, in total, we have 17 indices. Index 0 is the initial activation function. The odd indices are related to the activation function in the main path, and the even indices correspond to activation functions after the addition of the residual connection to the main path output. We select two scenarios. In the first scenario, denoted by Single Layer, we present the results when MeanSparse is applied to only the corresponding row-indexed activation function, while Cumulative depicts the result when MeanSparse is applied to the activation functions with indices less than or equal to the index represented in each row.
One shared remark in both Single Activation and Cumulative cases is that the starting layers are more effective in improving robustness. We can attribute this to the abstraction level in different layers of the ResNet18 model. In the case of Single Activation, when we apply the MeanSparse technique to larger indices, we observe that not only does the robust accuracy decrease, but the clean accuracy increases. This indicates that for even better performance, we can use different $\alpha$ values for activation functions based on their depth level. Note that for the current version of the MeanSparse technique, we use a shared $\alpha$ for all activation functions.
@@ A.6 Applying to Specific Activation Functions
Table 7: Clean and APGD accuracy of ResNet18 model with GELU activation function over CIFAR-10 test set after the application of MeanSparse technique for Single Activation and Cumulative cases. Single Activation represents the results when MeanSparse is applied to only the index activation function, and Cumulative Activation represents the results when MeanSparse is applied to the index activation function and all previous activation functions (for the based model the Clean and APGD accuracy are $84.59\%$ and $49.12\%$, respectively).
@@
Table 8: Clean and APGD accuracy of ResNet18 model with GELU activation function over CIFAR-10 test set before (Base denoted by $\alpha=0$ in the table) and after the application of MeanSparse technique for Main Path (application of MeanSparse to activation functions in the main path of residual block) and After Addition (application of MeanSparse to activation functions operating after addition of main path result and residual data) cases.

### v2

@@ 4.3 Insights from Ablation Study
We conducted several experiments to explore different aspects of MeanSparse integration into trained models. These experiments examine the effects of activation functions A.1, the type of adversarial attack (either $\ell_{\infty}$ or $\ell_{2}$) A.2, the robustness of MeanSparse integrated models against black-box attacks A.3, the effect of adaptive attack with full access to MeanSparse modules location and specifications (mean and variance vectors) A.4, the types of adversarial examples used during training with AT A.5, the effect of integrating a standard trained model with MeanSparse A.6, the impact of centering MeanSparse around the mean value A.7, the influence of attack power on MeanSparse integration performance A.8, the effects of integrating MeanSparse with different patterns, such as isolated integration with a single activation A.9, and the robustness improvement resulting from MeanSparse integration in terms of PGD-50 attack A.10. The main takeaways from our ablation study can be summarized as:
@@
  • By calculating the sparsity using other well-known adversarial attacks such as PGD-50 [29], we observe a similar improvement in robustness, consistent with the AA and APGD accuracy values.
  • Although MeanSparse masks the gradient, which can explain the robustness improvement against white-box attacks, it also demonstrates enhanced robustness when evaluated using black-box attacks.
  • Designing adaptive adversarial attacks for models integrated with MeanSparse can slightly reduce the robust accuracy. However, due to the significant improvement from MeanSparse integration, the adaptively attacked models still exhibit substantially higher robustness compared to the original base model (see Section A.4 for a detailed description of the adaptive attack).
@@ 4.4 Limitations
The MeanSparse technique proposed in this paper is a post-training method that is useful on top of adversarially trained models. Using MeanSparse for models trained in a standard way does not lead to robustness improvement as we expect based on our intuition introduced in this paper.
@@ 5 Conclusions
Adversarial training, a widely accepted defense against evasion attacks, attenuates non-robust feature extractors in a model. Although adversarial training (AT) is effective, this paper demonstrates that there is still an easily accessible capacity for attackers to exploit. We introduce the MeanSparse technique, which partially blocks this capacity in the model available to attackers and enhances adversarial robustness. This technique, easily integrable after training, improves robustness without compromising clean accuracy. By integrating this technique into trained models, we achieve a new record in adversarial robustness on CIFAR-10 for both $\ell_{\infty}$ and $\ell_{2}$ untargeted attacks, as well as on CIFAR-100 and ImageNet for $\ell_{\infty}$ untargeted attacks.
@@ Appendix A Ablation Experiment
We conducted several experiments to evaluate the effectiveness of MeanSparse in various scenarios. In the majority of experiments, we use APGD accuracy (accuracy after attacking the model with Auto-PGD using both Cross-entropy and Difference of Logits Ratio loss metrics) as the robustness metric [10]. During our simulations, we found APGD accuracy to be almost equal to AA accuracy, with the advantage that APGD can be measured faster.
@@ A.1 Activation Function
Table 2: Clean and APGD accuracy of ResNet-18 model with different activation functions over CIFAR-10 test set before (Base) and after integration with MeanSparse technique for different value of $\alpha$
@@ A.2 Threat Model
In this experiment, we focused on the effectiveness of MeanSparse against two well-studied threat models, $\ell_{\infty}$ and $\ell_{2}$. Until now, our focus was on the $\ell_{\infty}$ threat model, which generally results in lower robust accuracy [9]. On the other hand, $\ell_{2}$ threat models can focus on a specific part of the image, and this localization of adversarial perturbation may challenge the effectiveness of the MeanSparse technique. In this experiment, we compare the robustness against both $\ell_{\infty}$ and $\ell_{2}$ threat models. The model is ResNet-18 with the GELU activation function as in our Activation Function experiment in 2. Table 3 illustrates the clean and APGD accuracy for $\ell_{\infty}$ and $\ell_{2}$ threats. As the accuracy for the $\ell_{\infty}$ threat model increases with increasing $\alpha$, the accuracy for $\ell_{2}$ also increases, which depicts the effectiveness and generalization of MeanSparse across well-studied threat models. Additionally, the improvement compared to the base model at $\alpha=0.2$ is larger for the $\ell_{2}$ threat than for the $\ell_{\infty}$ threat.
Table 3: Clean and APGD accuracy of ResNet-18 model with GELU activation function over CIFAR-10 test set before (Base) and after integration with MeanSparse technique for different value of $\alpha$ and different threat models from RobustBesnch [10]
@@
A.3 Blackbox Attacks
The MeanSparse operation masks the gradient during backward propagation when the input $a^{\mathrm{in}}$ lies within $\alpha\sigma_{\mathrm{ch}}$ of the channel mean $\mu_{\mathrm{ch}}$ (see the bottom plot in the first column of Figure 1). An interesting direction to explore is the performance of black-box attacks that do not rely on gradients, both before and after integrating MeanSparse. For this experiment, we select the Square Attack [2]. Table 4 shows the clean and robust accuracy for a ResNet-18 model with GELU activation before (Base) and after integration with MeanSparse (robust accuracy is denoted by Square, representing the accuracy under Square Attack). The results indicate a similar improvement in robust accuracy on par with APGD, while clean accuracy remains largely unaffected across a wide range of $\alpha$ values, from $0.05$ to $0.2$.
Table 4: Clean and Square accuracy (Square accuracy in the Square column represents the accuracy after applying square attack [2]) of ResNet-18 model with GELU activation function over CIFAR-10 test set before (Base) and after integration with MeanSparse technique for different value of $\alpha$
$\alpha$ Clean Square
Base $84.59$ $57.37$
$0.05$ $\mathbf{84.63}$ $57.66$
$0.10$ $84.58$ $58.19$
$0.15$ $84.42$ $59.31$
$0.20$ $84.32$ $60.30$
$0.25$ $84.18$ $60.67$
$0.30$ $83.95$ $61.32$
$0.35$ $83.51$ $\mathbf{61.52}$
$0.40$ $82.85$ $61.49$
A.4 Adaptive Attack
In this experiment, we investigate the potential for the attacker to enhance the attack by utilizing the knowledge that MeanSparse is employed to robustify the model. We propose a formulation to adapt the original PGD attack, incorporating the integration of MeanSparse to create a more powerful attack. In this scenario, we assume the attacker has complete access to the locations, mean, and variance vectors of the integrated MeanSparse operator throughout the model. Let $\bm{x}$ represent the input image and $\{\bm{z}_{i}\}$ the corresponding representations at the input of the MeanSparse modules in the model. When a perturbation $\bm{\delta}$ is added to the input ($\bm{x}+\bm{\delta}$), it results in new representations $\{\bm{z}_{i}^{\delta}\}$. In the adaptive PGD attack, we solve the following optimization problem to determine the perturbation $\bm{\delta}$:
$\operatorname*{arg\,max}_{\bm{\delta}}\quad\mathrm{CE}(f(\bm{x}+\bm{\delta}),y)-\lambda\sum_{i}\|\bm{m}_{i}\odot(\bm{z}_{i}-\bm{z}_{i}^{\delta})\|_{p}^{q}$
$\text{subject to}\quad\|\bm{\delta}\|_{\infty}\leq\epsilon,\quad\bm{x}+\bm{\delta}\in[0,1]^{D}$ (9)
where $\mathrm{CE}(\cdot,\cdot)$ is the cross-entropy loss, $f$ is the model, $y$ is the true label, $\lambda$ is a hyper-parameter, $i$ indexes different MeanSparse modules in the model, $\odot$ is the element-wise multiplication operator, $\epsilon$ controls the perturbation imperceptibility via the $\ell_{\infty}$ norm, and $D$ is the dimension of the input image. $\bm{m}_{i}$ represents a mask with the same dimensions as $\bm{z}_{i}$, identifying the locations in $\bm{z}_{i}$ affected by the MeanSparse module. For a given $\bm{z}$, the corresponding $\bm{m}$ can be computed element-wise as:
$m_{j}=\begin{cases}1&\text{if~{}~{}}\mu_{\mathrm{ch}}-\alpha\sigma_{\mathrm{ch}}\leq z_{j}\leq\mu_{\mathrm{ch}}+\alpha\sigma_{\mathrm{ch}}\\ 0&\text{Otherwise}\end{cases},j=1,2,\ldots,K$
where $\mu_{\mathrm{ch}}$ and $\sigma_{\mathrm{ch}}$ are the statistics of the MeanSparse module, and $K$ represents the representation dimension. Thus, whenever a feature in the input of MeanSparse is affected by this module for the input image $\bm{x}$, the corresponding mask value becomes $1$; otherwise, it is 00. As a result, the second term in the objective function of the adaptive PGD attack equation 9 penalizes perturbations that lead to changes in the features affected by MeanSparse. In the ideal case, the perturbation is designed such that none of the MeanSparse modules can block its effect in their input (we know that the changes in the input of the $i$-th MeanSparse module are $\bm{z}_{i}-\bm{z}_{i}^{d}$). Thus, the attacker can bypass the MeanSparse, which we denote as the Adaptive-$(p,q)$ attack.
To test the adaptive attack, we solve the optimization problem equation 9 using gradient ascent with a step size of $\mu=\frac{1}{255}$, a number of iterations $N=50$, and $\epsilon=\frac{8}{255}$. When $\lambda=0$, problem equation 9 reduces to the well-known PGD-50 attack, while for larger values of $\lambda$, we obtain an adaptive attack. We conducted experiments using this adaptive attack on a ResNet-18 model with a GELU activation function, which was adversarially trained on the CIFAR-10 dataset. Table 5 compares the results of the PGD-50 attack with the Adaptive-$(1,1)$ and Adaptive-$(2,2)$ attacks before (Base) and after integration ($\alpha=0.20$ and $\alpha=0.35$) with MeanSparse. For the base model, all the attacks are equivalent, as $\bm{m}_{i}$ is a zero matrix for all MeanSparse modules. As $\alpha$ increases, the accuracy after applying our adaptive attack decreases compared to the PGD-50 attack, demonstrating the effectiveness of our adaptation. While the adapted attack is more powerful than the original PGD-50 attack, the improvement gap in robustness resulting from MeanSparse is still significant. For the base model, the robust accuracy is $51.56\%$, while the worst accuracies are $53.56\%$ and $55.26\%$ for $\alpha=0.20$ and $\alpha=0.35$, respectively, indicating a clear improvement in robustness, while the clean accuracy remains almost unchanged.
Table 5: Clean, PGD-50 and Adaptive Attack accuracy for a ResNet-18 model with GELU activation over CIFAR-10 dataset before (Base) and after integration with MeanSparse
$\alpha$ Clean PGD-50 Adaptive-$(1,1)$ Adaptive-$(2,2)$
Base $\mathbf{84.59}$ $51.56$ $51.56$ $51.56$
$0.20$ $84.32$ $53.89$ $53.56$ $53.58$
$0.35$ $83.51$ $\mathbf{55.50}$ $\mathbf{55.36}$ $\mathbf{55.26}$
A.5 Adversarial Training
@@
A.6 Standard Training
In the previous experiments, we explore the effect of two different adversarial training methods on performance improvement by MeanSparse. In this experiment, we test MeanSparse integration over a standard trained model. Table 7 represents the APGD accuracy over a ResNet-18 model with standard training for $\epsilon$ values starting from $1/255$ to $8/255$. The first column ($\alpha=0$) is the base model performance while the next two columns ($\alpha=0.2$ and $\alpha=0.35$) represent the models integrated with MeanSparse. In all cases, the attacker can fool the model with imperceptible perturbation ($\epsilon\leq 4/255$) which necessitates the integration of MeanSparse with adversarially trained models.
Table 7: APGD accuracy of standard trained ResNet-18 model with GELU activation function over CIFAR-10 test set before (Base denoted by $\alpha=0$ in the table) and after the integration of MeanSparse technique for different attack power $\epsilon$ and two different $\alpha$ values ($0.20$ and $0.35$). Clean accuracy is provided next to each $\alpha$ value.
$\alpha$
0 ($94.36$) 0.2 ($94.20$) 0.35 ($93.61$)
$\epsilon$ APGD APGD APGD
$1/255$ $49.09$ $52.06$ $56.36$
$2/255$ $4.86$ $7.69$ $14.52$
$3/255$ $0.07$ $0.24$ $1.33$
$4/255$ $0.0$ $0.01$ $0.03$
$5/255$ $0.0$ $0.0$ $0.01$
$6/255$ $0.0$ $0.0$ $0.01$
$7/255$ $0.0$ $0.0$ $0.01$
$8/255$ $0.0$ $0.0$ $0.01$
A.7 Feature Centerization Type
The MeanSparse technique applies sparsification over mean-centered features across each channel in the input tensor to the activation function. In this section, we inspect the importance of the reference for feature centralization before sparsification in our proposed technique. For this purpose, we compare three cases where the reference for centering features is the channel mean ($\bm{\mu}_{ch}$), zero, and the mean over all features in all channels ($\mu$). For a fair comparison across all cases, we select the sparsification threshold equal to $\alpha\times\sigma_{ch}$, where $\sigma_{ch}$ emphasizes the channel-wise calculation standard deviation. Table 8 represents the result for different options of reference. When using zero as the reference for feature centralization, the improvement in APGD is higher than using the mean as the reference, but on the other hand, the utilization of the model also decreases. One challenging problem is the high rate of clean accuracy reduction when using zero as the reference, which may cause a huge reduction if not selected properly. When using the mean of all channels as the reference, the results in terms of both clean and APGD accuracy metrics underperform compared to the case where the channel mean is used as the reference. The main takeaway from the table is the importance of using channel-wise statistics to maintain utility while improving robustness.
Table 8: Clean and APGD accuracy of ResNet-18 model with GELU activation function over CIFAR-10 test set before (Base) and after the application of MeanSparse technique for different references for feature centralization
@@ A.8 Attack Power
Table 9: APGD accuracy of ResNet-18 model with GELU activation function over CIFAR-10 test set before (Base denoted by $\alpha=0$ in the table) and after the application of MeanSparse technique for different attack power $\epsilon$ and two different $\alpha$ values ($0.20$ and $0.35$). Clean accuracy is provided next to each $\alpha$ value.
@@
A.9 Applying to Specific Activation Functions
The MeanSparse technique is designed to sparsify mean-centered inputs to all activation functions in a model. In this experiment, we will focus on applying sparsification over a selected subset of activation functions. For this purpose, we index the activation functions in the ResNet-18 model. This model has 8 residual blocks, each with 2 activation functions: one in the Main Path and one after the result of the main path is added to the residual connection, totaling 16 activation functions. Additionally, this architecture has one activation function before the residual block. So, in total, we have 17 indices. Index 0 is the initial activation function. The odd indices are related to the activation function in the main path, and the even indices correspond to activation functions after the addition of the residual connection to the main path output. We select two scenarios. In the first scenario, denoted by Single Activation, we present the results when MeanSparse is applied to only the corresponding row-indexed activation function, while Cumulative depicts the result when MeanSparse is applied to the activation functions with indices less than or equal to the index represented in each row.
Table 10 represents the results for both Single Activation and Cumulative scenarios. One shared remark is that the starting layers are more effective in improving robustness. We can attribute this to the abstraction level in different layers of the ResNet-18 model. In the case of Single Activation, when we apply the MeanSparse technique to larger indices, we observe that not only does the robust accuracy decrease, but the clean accuracy increases. This indicates that for even better performance, we can use different $\alpha$ values for activation functions based on their depth level. Note that for the current version of the MeanSparse technique, we use a shared $\alpha$ for all activation functions.
@@ A.9 Applying to Specific Activation Functions
Table 10: Clean and APGD accuracy of ResNet-18 model with GELU activation function over CIFAR-10 test set after the application of MeanSparse technique for Single Activation and Cumulative cases. Single Activation represents the results when MeanSparse is applied to only the index activation function, and Cumulative Activation represents the results when MeanSparse is applied to the index activation function and all previous activation functions (for the based model the Clean and APGD accuracy are $84.59\%$ and $49.12\%$, respectively).
@@
Table 11: Clean and APGD accuracy of ResNet-18 model with GELU activation function over CIFAR-10 test set before (Base denoted by $\alpha=0$ in the table) and after the application of MeanSparse technique for Main Path (application of MeanSparse to activation functions in the main path of residual block) and After Addition (application of MeanSparse to activation functions operating after addition of main path result and residual data) cases.
@@
A.10 Attack Type
Figure LABEL:fig:spaf demonstrates the improvement in robustness of leading models in RobustBench in terms of AutoAttack accuracy. In this experiment, we use PGD-50 [29] for robustness measurement. Table 12 represents the Clean and PGD-50 accuracy values for different models before (denoted by Base) and after integration with MeanSparse (denoted by MeanSparse) over CIFAR-10, CIFAR-100 and ImageNet datasets. The results again confirm the improvements in Figure LABEL:fig:spaf as t integration with MeanSparse increases the PGD-50 accuracy in all models.
Table 12: Clean and PGD-50 [29] accuracy of various RobustBench-ranked models [9], both before (Base) and after integration with the MeanSparse technique (models tested here includes WideResNet-94-16 [3] and RaWideResNet-70-16 [34] over CIFAR-10 dataset, WideResNet-70-16 [48] over CIFAR-100 dataset and Swin-L [27] and ConvNeXt-L [26] over ImageNet dataset
CIFAR-10 CIFAR-100 ImageNet
WideResNet-94-16 RaWideResNet-70-16 WideResNet-70-16 Swin-L ConvNeXt-L
Clean PGD-50 Clean PGD-50 Clean PGD-50 Clean PGD-50 Clean PGD-50
Base $\mathbf{93.68}$ $76.36$ $\mathbf{93.27}$ $74.01$ $\mathbf{75.23}$ $48.35$ $\mathbf{78.92}$ $61.42$ $\mathbf{78.02}$ $60.62$
MeanSparse $93.63$ $\mathbf{78.21}$ $93.24$ $\mathbf{75.17}$ $75.17$ $\mathbf{50.64}$ $78.88$ $\mathbf{61.78}$ $77.96$ $\mathbf{61.78}$